A manufacturer is making a cylindrical can that will hold and dispense flower seeds through small holes in the top of the can. The manufacturer wants the can to have a volume of 42 cubic inches and be 6 inches tall. What should the diameter of the can be? V=πr. Round your answer to the nearest inch.
If possible please show me a step by step process.

Answer by lwsshak3(11628)

A manufacturer is making a cylindrical can that will hold and dispense flower seeds through small holes in the top of the can. The manufacturer wants the can to have a volume of 42 cubic inches and be 6 inches tall. What should the diameter of the can be? V=πr. Round your answer to the nearest inch.
If possible please show me a step by step process.
***
let h=height=6 inches
let d=diameter
Area of circular ends=πr^2 or πd^2/4
volume=area*height=(πd^2/4)*6
6πd^2/4=42
6πd^2=4*42=168
d^2=168/6π=28/π
d=2.99
What should the diameter of the can be? 3 inches